A loan of$2400 is to be repaid by 20 equal annual instalments. The rate of interest for the transaction is 10%per annum. Find the amount of each annual repayment, assuming that payments are made
(a) in arrears
(b) in advance
1
Expert's answer
2020-10-05T16:57:53-0400

(a) in arears


"\frac{r}{100}\times 2400"


"\frac{10}{100}\times2400"


"0.1\times2400"


"=\$240"


(b) in advance


"v=\frac{1}{1+i}"


"v=\frac{1}{1+0.07}"


"v=\frac{1}{1.01}"


"v=0.909090"


"2400\times0.909090"


"=2181.816"

"2400-2181.816"


"=\$218.184"